OpenStudy (anonymous):

MEDAL!? Find the intersection points, if any, for each system of equations: (x^2/4)+y^2=1 y=x+1

OpenStudy (marigirl):

The first equation is an ellipse. so both the equations can be graphed

OpenStudy (marigirl):

So then the graphs look like this

OpenStudy (marigirl):

then we can find the points of intersections: (0,1) and (-1.6,-0.6)

OpenStudy (marigirl):

that was done graphically.. Does your answer need to be done algebraically? Let me know :)

OpenStudy (anonymous):

Is that what I would write as my answer then? If it is not too much trouble could you write about the process?

OpenStudy (marigirl):

What I did above was to graph it using a graphics calculator and find the points of intersection through that. If you were to solve it algebraically we can use the substitution method: (x^2/4)+y^2=1 and y=x+1 using substition \[\frac{ x^2 }{ 4 }+(x+1)^2=1\] Then we solve for x and the two x solutions that you will obtain will be x=0 and x=-1.6. Then you take those two values and plug it into any of the two original equations to obtain the coordinates (0,1) and (-1.6,-0.6) So the line y=x+1 crosses the ellipse at two places.

OpenStudy (anonymous):

Oh ok! What about x^2+y^2-16y+39=0 y^2-x^2-9=0

OpenStudy (marigirl):

ellipse and hyperbola..

OpenStudy (anonymous):

ANy intersection points?

OpenStudy (marigirl):

Make y the subject of the second equation and then substitute into the first..give me some time for the intersections lol

OpenStudy (marigirl):

THere are points that intersect but it is not showing up on my calc :(

OpenStudy (anonymous):

try desmos.com/calculator

OpenStudy (marigirl):

Ill tell you what I have so far to graph this The ellipse: \[x=\sqrt{-y^2+16y-39}\]

OpenStudy (marigirl):

\[x=-\sqrt{-y^2+16y-39}\]

OpenStudy (marigirl):

The Hyperbola \[y=\sqrt{x ^2+9}\] \[y=-\sqrt{x ^2+9}\]

OpenStudy (marigirl):

I gotta go now, sorry i can come back later and solve it by hand .i have a feeling the ans could be irrational

OpenStudy (marigirl):

and plz give me a medal for the first one :)

OpenStudy (anonymous):

Please come back later! I will give you medals if you can help me solve just this last one :)

OpenStudy (marigirl):

I got it :D :D

OpenStudy (marigirl):

I first took the circle equation x^2+y^2-16y+39=0 and make it as \[y=8-\sqrt{25-x ^{2}}\] then Intersect with \[y=\sqrt{x^2+9}\]

OpenStudy (marigirl):

so it looks like this

OpenStudy (marigirl):

so the 3 points of intersection are (-4,5) (0,3) and (4,5)